OpenStudy (kimberly_pr):

Which of the following is the solution to the equation 25(z + 4) = 125? z = 5.5 z = 3.5 z = −2.5 z = −0.5

OpenStudy (kimberly_pr):

The equation is 25^(z+4)=125

OpenStudy (kimberly_pr):

@naate

OpenStudy (kimberly_pr):

@welshfella @ganeshie8

OpenStudy (naate):

lemme give it a shot lol

OpenStudy (welshfella):

take logs of both sides

OpenStudy (welshfella):

and use the following rule of logs log a^b = b log a

OpenStudy (naate):

z is 1

OpenStudy (kimberly_pr):

How do I turn that into a log? :/

OpenStudy (naate):

i found the answer, but its not on there, unless im trippin

OpenStudy (kimberly_pr):

And yeah, that's what I got but it's not an answer choice, your supposed to solve it using a logarithmic function but I SUCK at that @naate

OpenStudy (welshfella):

log 25^(z + 4) = log 125 (z + 4) log 25 = log 125

OpenStudy (naate):

ok lol i guess i do too @Kimberly_PR

OpenStudy (welshfella):

now divide both sides by log 25

OpenStudy (kimberly_pr):

Ok... So you would get (z+4) = 1.5 Am I right so far?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

now its easy

OpenStudy (kimberly_pr):

Now I just need to solve for z like a regular equation... SO I got z= -2.5 Is that correct?

OpenStudy (welshfella):

yea

OpenStudy (kimberly_pr):

Ok, thank you soooo much! @welshfella And @naate thank you too for trying :) lol

OpenStudy (welshfella):

yw

OpenStudy (naate):

lol no prob anytime @Kimberly_PR

OpenStudy (kimberly_pr):

I gave @welshfella the medal for best response but I still fanned you so thanks again, lol I guess we both learned something new today @naate

OpenStudy (naate):

thnx! @Kimberly_PR